OpenStudy (anonymous):

Six differently colored balls, (red, blue, green, orange, purple, and write) are placed in a basket. Without looking, three balls are removed. What is the total number of combinations that include a red ball?

OpenStudy (dumbcow):

i've seen this exact question on a practice HS AIMS test...are you from AZ :) anyway, since you know 1 of the 3 is a red ball, the problem is simply how many ways can you pick 2 balls out of the remaining 5? 5 choose 2 \[_{5}C_{2} = \frac{5!}{2!3!} = \frac{5*4*3!}{2*3!} = \frac{20}{2}=10\]